OpenStudy (anonymous):

Score Number of Students 0 2 1 1 2 3 3 6 4 2 5 3 6 2 Based on the table, what is the mean hockey score?

OpenStudy (anonymous):

\[\frac{ total number }{number of scores ? }\]

OpenStudy (anonymous):

Oops, does 2,1,3,6,2,3,2 in the number of student category?

OpenStudy (anonymous):

belong in the #* ??

OpenStudy (anonymous):

yes @Shalante

OpenStudy (anonymous):

total score=(number # of students with score 0)(0)+(number # of students with score 1)(1)+(# of students with score 2)(2)+(# of students with score 3)(3)+.....

OpenStudy (anonymous):

total score/number of students for mean

OpenStudy (anonymous):

mean=total score/number of students.

OpenStudy (anonymous):

@Shalante

OpenStudy (anonymous):

i got 2.7

OpenStudy (anonymous):

let me see

OpenStudy (anonymous):

Nope

OpenStudy (anonymous):

What was the total score for yours?

OpenStudy (anonymous):

um 19

OpenStudy (anonymous):

No. I mentioned earlier you were supposed to multiply the amount of student by the scores. Mean means average. Total of all scores by including every student and scores/total number of student This is how you find the average score.

OpenStudy (anonymous):

114

OpenStudy (anonymous):

@Shalante

OpenStudy (anonymous):

No. Just pay attention. 1st part: Total score Two students with score of 0 2x0=0 One student with score of 1 1x1=1 Three students with score of 2 3x2=6 Six students with score of 3 6x3=18 Two students with a score of 4 2x4=8 Three students with a score of 5 3x5=15 Two students with a score of 6 2x6=12 Add all the scores up 0+1+6+18+8+15+12=60 Total score is 60 Number of students 2+1+3+6++2+3+2=17 Mean=total scores/number of results or number of students Mean=60/17 =3.52 Understand now?

OpenStudy (anonymous):

these are the choices though 2.7 2.9 3.2 5.2

OpenStudy (anonymous):

@Shalante

OpenStudy (anonymous):

oops 60/19

OpenStudy (anonymous):

number of student is 19 lol. my mistake

OpenStudy (anonymous):

then you round it.

OpenStudy (anonymous):

so 3.2 ? @Shalante

OpenStudy (anonymous):

@Shalante

OpenStudy (anonymous):

yes.